# Learning WENO for entropy stable schemes to solve conservation laws

### v1

@@ 5.1 Reconstruction Accuracy
Figure 4: Reconstruction of a simulated discontinuity, which is not well-resolved, using: (a) ENO3, (b) SP-WENO, (c) SP-WENOc, (d) DSP-WENO.
5.2 Linear Advection
Consider the one-dimensional scalar linear advection equation
$u_{t}+cu_{x}=0.$
For the 1D scalar case, the finite difference TeCNO4 numerical flux (13) simplifies to
$f_{i+\frac{1}{2}}=\tilde{f}^{2p}_{i+\frac{1}{2}}-a_{i+\frac{1}{2}}\left(v_{i+\frac{1}{2}}^{+}-v_{i+\frac{1}{2}}^{-}\right),$ (40)
where $v:=\eta^{\prime}(u)=u$ is the entropy variable corresponding to the square entropy $\eta(u)=\frac{u^{2}}{2}$. The entropy conservative component $\tilde{f}^{2p}$ of the flux is constructed using (7) with the following second-order entropy conservative flux
$\tilde{f}_{i+\frac{1}{2}}=c\frac{(u_{i}+u_{i+1})}{2}.$
Additionally, $a_{i+\frac{1}{2}}=|c|$ for the linear advection model. We set the convective velocity $c=1$ for all test cases.
5.2.1 Smooth Initial Condition
We consider two test cases with smooth initial conditions on $[-\pi,\pi]$ with periodic boundary conditions.
Table 4 shows the $L_{h}^{1}$ errors with various reconstructions for both of these test cases. ENO3 works well for Test 1 and achieves third-order convergence with marginally smaller absolute errors when compared to the results of SP-WENO, SP-WENOc, and DSP-WENO. However, the convergence rate with ENO3 deteriorates in Test 2. It has been observed in [48] that ENO3 performs poorly for this test case with the MUSCL scheme due to linear instabilities, and we observe the same issues with ENO3 in the TeCNO framework (also seen in [16]). The other reconstruction methods are stable, as SP-WENO, SP-WENOc, and DSP-WENO do not suffer from this issue. All three reconstruction methods achieve third-order convergence, though the errors of DSP-WENO are about an order of magnitude larger than those of SP-WENO and SP-WENOc. Moreover, SP-WENO and SP-WENOc exhibit super-convergence. This can be attributed to the reconstructed jump often being zero, which effectively disables the local diffusion and results in the local numerical flux entirely consisting of the fourth-order entropy conservative flux.
DSP-WENO is much more dissipative, so in these test cases and in general for smooth regions, we observe larger errors (though the third-order convergence is still maintained). However, this increase in diffusion leads to improved performance over SP-WENO and SP-WENOc when the solution is not smooth (see the results of Sections 5.2.2, 5.3, and 5.4).
N ENO3 SP-WENO SP-WENOc DSP-WENO
$L_{h}^{1}$ $L_{h}^{1}$ $L_{h}^{1}$ $L_{h}^{1}$
Error Rate Error Rate Error Rate Error Rate
100 3.23e-5 - 6.90e-5 - 6.80e-5 - 1.66e-4 -
200 4.04e-6 3.00 7.65e-6 3.17 7.48e-6 3.18 3.58e-5 2.21
Test 1 400 5.05e-7 3.00 8.29e-7 3.20 8.17e-7 3.20 4.57e-6 2.97
600 1.50e-7 3.00 2.26e-7 3.20 2.23e-7 3.20 1.35e-6 3.02
800 6.31e-8 3.00 8.72e-8 3.31 8.60e-8 3.31 5.72e-7 2.97
1000 3.23e-8 3.00 4.21e-8 3.27 4.15e-8 3.26 2.95e-7 2.97
100 1.48e-3 - 1.52e-3 - 1.46e-3 - 1.87e-3 -
200 1.98e-4 2.91 1.68e-4 3.18 1.68e-4 3.12 2.61e-3 2.84
Test 2 400 2.58e-5 2.94 1.79e-5 3.23 1.78e-5 3.23 3.35e-5 2.96
600 8.25e-6 2.81 4.69e-6 3.31 4.70e-6 3.29 9.59e-6 3.08
800 4.64e-6 2.00 1.81e-6 3.31 1.80e-6 3.33 3.93e-6 3.10
1000 3.46e-6 1.31 8.64e-7 3.32 8.61e-7 3.31 2.03e-6 2.96
Table 4: Linear advection Tests 1 and 2 with TeCNO4 flux.
5.2.2 Discontinuous Initial Condition
Test 3 The domain is $[0,1.4]$, final time is $T=1.4$, and CFL = 0.2, with initial profile
$u_{0}(x)=\begin{cases}10(x-0.2)&\text{if }0.2<x\leq 0.3\\ 10(0.4-x)&\text{if }0.3<x\leq 0.4\\ 1&\text{if }0.6<x\leq 0.8\\ 100(x-1)(1.2-x)&\text{if }1.0<x\leq 1.2\\ 0&\text{otherwise}\end{cases},$
and periodic boundary conditions. Note that the initial profile is a composition of shapes with different orders of regularity. The results on a mesh with 100 cells for various reconstruction methods are shown in Figure 5. ENO2 leads to the most dissipative approximation on this mesh. Overall, ENO3 yields the best performance, as its solution is the closest to the exact solution while avoiding any spurious oscillatory behavior. The solution obtained with DSP-WENO appears to be marginally more dissipative than ENO3, but without exhibiting the large undershoots/overshoots observed when using SP-WENO or SP-WENOc.
Figure 5: Linear advection Test 3: Solution with TeCNO4 at time $T=1.4$. (a) Comparison of different reconstruction methods for $x\in[0,1.4]$, (b) Zoomed in comparison for $x\in[0.4,1]$.
5.3 Burgers’ Equation
We now consider the Burgers’ Equation
$u_{t}+\left(\frac{u^{2}}{2}\right)_{x}=0.$
As with linear advection, we use (40) with entropy variables $v=u$ corresponding to the square entropy $\eta(u)=\frac{u^{2}}{2}$. The high-order entropy conservative flux (7) makes use of the following second-order entropy conservative flux
$\tilde{f}_{i+\frac{1}{2}}=\frac{(u_{i}^{2}+u_{i+1}^{2}+u_{i}u_{i+1})}{6}.$
Additionally, we take $a_{i+\frac{1}{2}}=\frac{|u_{i}|+|u_{i+1}|}{2}$.
Test 1: The domain is $[-1,1]$, final time is $T=0.5$, and CFL = 0.4, with initial profile
$u_{0}(x)=\begin{cases}3&\text{if }x<0\\ -1&\text{if }x>0\end{cases},$
and Neumann boundary conditions. The mesh consists of 100 cells. Figure 6 shows that all the solutions experience oscillations. This is a fundamental problem with high-order TeCNO schemes, so there often is no way to completely eliminate the overshooting behavior without introducing additional diffusion. However, we observe that DSP-WENO performs better in mitigating the oscillations leading up to the overshoot.
Remark 10.
While omitted from Figure 6, we found that these oscillations cannot be avoided even with the most restrictive linear reconstruction using the minmod limiter, which also satisfies the sign property [14].
Figure 6: Burgers’ Test 1: Solution with TeCNO4 at time $T=0.5$. (a) Comparison of different reconstruction methods for $x\in[-1,1]$, (b) Zoomed in comparison for $x\in[0.3,0.5]$.
$u_{0}(x)=\begin{cases}3&\text{if }-1\leq x<-0.5\\ 1&\text{if }-0.5\leq x<0\\ 3&\text{if }0\leq x<0.5\\ 2&\text{if }0.5\leq x<1\\ \sin{(\pi x)}&\text{if }|x|>1\end{cases},$
and periodic boundary conditions. This initial condition is taken from [45] and contains both smooth and discontinuous features. When solved on a mesh with 400 cells, Figure 7 shows that SP-WENO and SP-WENOc perform poorly near the shocks, especially at $x=1.3$ where the overshoots are relatively large. While the oscillations are still present with DSP-WENO, they are significantly mitigated when compared to the other SP-WENO methods.
Figure 7: Burgers’ Test 2: Solution with TeCNO4 at time $T=0.4$. Comparison of different reconstruction methods for various domains: (a) $x\in[-4,4]$, (b) $x\in[0.86,1.65]$.
5.4 Euler Equations
We now consider the Euler equations described by
where $\rho$, $\textbf{u}=(\text{u}_{1},\text{u}_{2},\text{u}_{3})^{\top}$, and $p$ represent the fluid density, velocity, and pressure, respectively. Note that u is distinct from ${\bm{u}}$, with the latter representing the vector of conserved variables. $E$ is the total energy per unit volume given by
$E=\rho\left(\frac{|\textbf{u}|^{2}}{2}+e\right).$
where $e$ is the specific internal energy given by a caloric equation of state, $e=e(\rho,p)$. Here, we take the equation of state to be
$e=\frac{p}{(\gamma-1)\rho},$
where $\gamma=c_{p}/c_{v}$ denotes the ratio of specific heats. In all the test cases, we set $\gamma=1.4$.
For the Euler equations, a popular choice for the entropy-entropy flux pair [20] is
$\eta=-\frac{\rho s}{\gamma-1},\hskip 8.53581pt\bm{q}=[q_{1},q_{2},q_{3}]=-\frac{\rho\textbf{u}^{\top}s}{\gamma-1},$ (41)
where $s=\ln{(p)}-\gamma\ln{(\rho)}$ is associated with the thermodynamic specific entropy. The corresponding vector of entropy variables is given by
[…]
where $\beta=\rho/(2p)$.
In the TeCNO scheme, we use the second-order kinetic energy preserving and entropy conservative (KEPEC) flux [5] and (7) to construct the fourth-order entropy conservative flux. In 1D, the KEPEC flux is given as:
where $\tilde{p}=\overline{\rho}/(2\overline{\beta})$ and $\beta=\rho/(2p)$. $\hat{\rho}$, $\hat{\beta}$ represent the logarithmic averages (see [22, 5]) of the respective positive quantities.
For the diffusion term in 1D, we use the Roe-type diffusion operator
$\mathbf{D}_{i+\frac{1}{2}}=\tilde{\mathbf{R}}_{i+\frac{1}{2}}\mathbf{\Lambda}_{i+\frac{1}{2}}\tilde{\mathbf{R}}_{i+\frac{1}{2}}^{\top},$ (44)
where $\mathbf{R}_{i+\frac{1}{2}}$ consists of the eigenvectors of the flux Jacobian $\partial_{{\bm{u}}}{\bm{f}}$ and
$\mathbf{\Lambda}_{i+\frac{1}{2}}=\operatorname{diag}{\left(|u-a|,|u|,|u+a|\right)}_{i+\frac{1}{2}}.$ (45)
In (45), $a=\sqrt{\gamma p/\rho}$ is the speed of sound in air. The matrices $\mathbf{D}_{i+\frac{1}{2}}$ and $\mathbf{\Lambda}_{i+\frac{1}{2}}$ are evaluated at some averaged state. For additional details on the diffusion operator, and the flux formulation in higher dimensions, we refer the readers to [5, 44].
5.4.1 1D Euler Equations
We begin our study of the Euler equations in one dimension.
Modified Sod Shock Tube: This is a shock tube problem [54] solved on the domain is $[0,1]$ with initial profile
$(\rho,\text{u},p)=\begin{cases}(1,0.75,1)&\text{if }x<0.3\\ (0.125,0,0.1)&\text{if }x\geq 0.3\end{cases}.$
It is solved on a domain with 400 cells with Neumann boundary conditions, up to a final time of $T=0.2$ with CFL=0.4. Figure 8 shows that SP-WENO and SP-WENOc exhibit significant overshoots, especially near the shock at $x\approx 0.72$. DSP-WENO markedly improves on the performance of both methods by featuring only very minor overshoots in comparison. Further, the accuracy in shock-capturing is comparable to ENO3, with a sharper resolution of the contact discontinuity at $x\approx 0.57$.
Figure 8: Modified Sod Shock Tube Test: Solution with TeCNO4 at time $T=0.2$ for $\rho$, u, and $p$. Comparison of different reconstruction methods for various domains: (a) $x\in[0,1]$ for $\rho$, u, and $p$, (b) $\rho$: $x\in[0.5,0.8]$, u: $x\in[0.6,0.8]$, $p$: $x\in[0.6,0.8]$.
Shu-Osher Test: This test case, proposed in [50], contains the interaction of an oscillatory smooth wave and a right-moving shock. The domain is $[-5,5]$, final time is $T=1.8$, and CFL = 0.4, with initial profile
$(\rho,\text{u},p)=\begin{cases}(3.857143,2.629369,10.33333)&\text{if }x<-4\\ (1+0.2\sin{(5x)},0,1)&\text{if }x\geq-4\end{cases},$
and Neumann boundary conditions. The mesh consists of 400 cells. This was one of the test cases presented in [44] where SP-WENOc significantly mitigates the overshoots near the shock as compared to SP-WENO, which is also what we observe in Figure 9. DSP-WENO is clearly the most dissipative of the methods (although comparable to ENO3) when focusing in the regions with smooth physical high-frequency oscillations. However, there is essentially no overshoot near the shock with DSP-WENO, which is a large improvement over SP-WENO and SP-WENOc.
Figure 9: Shu-Osher Test: Solution with TeCNO4 at time $T=1.8$ for $\rho$, u, and $p$. Comparison of different reconstruction methods for various domains: (a) $x\in[-5,5]$ for $\rho$, u, and $p$, (b) $\rho$: $x\in[0,2.5]$, u: $x\in[-4,3]$, $p$: $x\in[-4,3]$.
Lax Test: For the Lax shock tube problem [29], the domain is $[-5,5]$, final time is $T=1.3$, and CFL = 0.4, with initial profile
$(\rho,\text{u},p)=\begin{cases}(0.445,0.698,3.528)&\text{if }x<0\\ (0.5,0,0.571)&\text{if }x\geq 0\end{cases},$
and Neumann boundary conditions. The mesh consists of 200 cells. Figure 10 shows that SP-WENO and SP-WENOc exhibit significant overshoots near the shocks. The solution obtained with DSP-WENO greatly minimizes the oscillatory behavior, similar to the previous test cases.
Figure 10: Lax Shock Tube Test: Solution with TeCNO4 at time $T=1.3$ for $\rho$, u, and $p$. Comparison of different reconstruction methods for various domains: (a) $x\in[-5,5]$ for $\rho$, u, and $p$, (b) $\rho$: $x\in[1.75,3.75]$, u: $x\in[-2.25,3.75]$, $p$: $x\in[-2.6,3.75]$.
Double Rarefaction: For this problem consisting of two rarefaction waves [34], the domain is $[-1,1]$, final time is $T=0.6$, and CFL = 0.4, with initial profile
$(\rho,\text{u},p)=\begin{cases}(7,-1,0.2)&\text{if }x<0\\ (7,1,0.2)&\text{if }x\geq 0\end{cases},$
and Neumann boundary conditions. This can be a challenging test case as the solution is close to vacuum near $x=0$. We solve the problem on a mesh consisting of 128 cells. With the absence of a discontinuity (except in the initial condition), the solutions obtained using ENO3, SP-WENOc, and DSP-WENO are very similar, as shown in Figure 11. Note that SP-WENO is omitted because the solver blows up due to the initial condition, even with lowering the CFL.
Figure 11: Double Rarefaction Test: Solution with TeCNO4 at time $T=0.6$ for $\rho$, u, and $p$. Comparison of different reconstruction methods for $x\in[-1,1]$ for $\rho$, u, and $p$.
5.4.2 2D Euler Equations
We continue our study of the Euler equations by considering several test cases in two dimensions.
Isentropic Vortex: In this problem, we consider the advection of a smooth isentropic vortex. This results in an entirely smooth solution, so we perform a mesh-refinement study here. The domain is $[-5,5]\times[-5,5]$, final time is $T=10$, and CFL = 0.5. The initial conditions are
$\text{u}_{1}=1-\frac{\Gamma}{2\pi}y\exp{\left(\frac{1-r^{2}}{2}\right)},\hskip 8.53581pt\text{u}_{2}=\frac{\Gamma}{2\pi}x\exp{\left(\frac{1-r^{2}}{2}\right)},\hskip 8.53581pt\mathcal{T}=1-\frac{(\gamma-1)\Gamma^{2}}{8\gamma\pi^{2}}\exp{(1-r^{2})},$
where $\mathcal{T}$ is the fluid temperature field, $r^{2}=x^{2}+y^{2}$ and $\Gamma=5$ (the vortex strength). Further, we have $\rho=\mathcal{T}^{1/(\gamma-1)}$ and $p=\mathcal{T}^{\gamma/(\gamma-1)}$ due to the isentropic conditions. Assuming periodic boundary condition, the vortex moves horizontal with unit velocity and completes one full periodic cycle at the final time. Table 5 shows $L_{h}^{1}$ errors with various reconstructions. We observe that ENO3 is unable to achieve third-order convergence, which we hypothesis is due to the linear instabilities with ENO [48]. The SP-WENO variants lead to third-order accuracy, with the results using DSP-WENO being more dissipative as compared to SP-WENO and SP-WENOc. We reiterate that this behavior is due to SP-WENO and SP-WENOc having zero reconstructed jumps in larger regions of the domain.
@@ 5.4 Euler Equations
50 1.17e-1 - 7.56e-2 - 7.51e-2 - 1.64e-1 -
100 1.75e-2 2.74 6.62e-3 3.51 6.73e-3 3.48 1.91e-2 3.10
Density 150 7.13e-3 2.22 1.37e-3 3.88 1.40e-3 3.87 5.25e-3 3.19
200 3.60e-3 2.38 4.59e-4 3.80 4.69e-4 3.80 2.14e-3 3.12
300 1.48e-3 2.19 1.01e-4 3.73 1.03e-4 3.74 6.16e-4 3.07
400 7.86e-4 2.20 3.57e-5 3.61 3.64e-5 3.61 2.50e-4 3.14
50 2.14e-1 - 1.80e-1 - 1.80e-1 - 3.04e-1 -
100 3.51e-2 2.61 2.10e-2 3.10 2.12e-2 3.09 4.39e-2 2.79
x-Velocity 150 1.30e-2 2.45 5.19e-3 3.44 5.26e-3 3.44 1.29e-2 3.03
200 6.06e-3 2.66 1.86e-3 3.58 1.87e-3 3.59 5.25e-3 3.12
300 2.27e-3 2.43 4.73e-4 3.37 4.75e-4 3.38 1.46e-3 3.15
400 1.17e-3 2.28 1.77e-4 3.41 1.79e-4 3.40 5.91e-4 3.15
50 1.71e-1 - 1.06e-1 - 1.07e-1 - 2.46e-1 -
100 2.52e-2 2.77 9.52e-3 3.48 9.75e-3 3.45 2.87e-2 3.10
Pressure 150 1.04e-2 2.18 1.99e-3 3.87 2.05e-3 3.85 7.81e-3 3.21
200 5.22e-3 2.40 6.60e-4 3.83 6.81e-4 3.82 3.19e-3 3.11
300 2.13e-3 2.21 1.43e-4 3.77 1.47e-4 3.77 9.13e-4 3.09
400 1.12e-3 2.23 5.02e-5 3.65 5.15e-5 3.66 3.70e-4 3.14
Table 5: $L_{h}^{1}$ error for advecting isentropic vortex with TeCNO4 flux.
2D Riemann problem (configuration 4): We consider a two-dimensional Riemann problem for the Euler equations [28] whose initial conditions are
$(\rho,\text{u}_{1},\text{u}_{2},p)=\begin{cases}(1.1,0,0,1.1)&\text{if }x>0.5,y>0.5\\ (0.5065,0.8939,0,0.35)&\text{if }x\leq 0.5,y>0.5\\ (1.1,0.8939,0.8939,1.1)&\text{if }x\leq 0.5,y\leq 0.5\\ (0.5065,0,0.8939,0.35)&\text{if }x>0.5,y\leq 0.5\end{cases}.$
When solved on the domain $[0,1]\times[0,1]$ with Neumann boundary conditions, the solution evolves into four interacting shock waves. We solve the problem on a mesh consisting of 400 $\times$ 400 cells until a final time $T=0.23$ with CFL = 0.5. The solutions depicted in Figure 12 seem qualitatively similar with all methods. However, if we notice the maximum solution values (listed under the colorbars), we can observe that it is the largest with SP-WENO, followed by SP-WENOc. Further, the maximum values with DSP-WENO are much closer to ENO3. We can attribute the variation of these values to overshoots near the shocks, which is captured more clearly in Figure 13 which shows one-dimensional slices of the solution at $x=0.5$ and $x=0.9$. Clearly, SP-WENO and SP-WENOc exhibit large overshoots, while DSP-WENO mitigates this behavior and is closer to ENO3. The reference solution is obtained using ENO3 by solving the problem on a finer $1200\times 1200$ mesh.
@@
Figure 12: 2D Riemann problem (configuration 4): Density profiles at time $T=0.23$. Comparison of TeCNO4 with different reconstruction methods: (a) ENO3, (b) SP-WENO, (c) SP-WENOc, (d) DSP-WENO.
Figure 13: 2D Riemann problem (configuration 4): One-dimensional slices of the solution using each reconstruction methods: (a) $x=0.9$, (b) $x=0.5$.
2D Riemann problem (configuration 12): We consider another two-dimensional Riemann problem for the Euler equations whose initial conditions on the domain $[0,1]\times[0,1]$ are given by
$(\rho,\text{u}_{1},\text{u}_{2},p)=\begin{cases}(0.5313,0,0,0.4)&\text{if }x>0.5,y>0.5\\ (1,0.7276,0,1)&\text{if }x\leq 0.5,y>0.5\\ (0.8,0,0,1)&\text{if }x\leq 0.5,y\leq 0.5\\ (1,0,0.7276,1)&\text{if }x>0.5,y\leq 0.5\end{cases}.$
In this case, the evolved solution (with Neumann boundaries) comprises two shock waves and two contact discontinuities. The problem is solved on a mesh consisting of 400 $\times$ 400 cells until a final time $T=0.25$ with CFL = 0.5. As shown in Figure 14, all methods sharply resolve the shocks and contact waves. However, SP-WENO and SP-WENOc lead to significant overshoots, which can once again be observed by looking at the maximum solution values and the one-dimensional slices shown in Figure 15. On the other hand, the overshoots are generally smaller with DSP-WENO and comparable to ENO3. The reference solution is once again generated using ENO3 by solving the problem on $1200\times 1200$ mesh.
@@
Figure 14: 2D Riemann problem (configuration 12): Density profiles at time $T=0.25$. Comparison of TeCNO4 with different reconstruction methods: (a) ENO3, (b) SP-WENO, (c) SP-WENOc, (d) DSP-WENO.
Figure 15: 2D Riemann problem (configuration 12): One-dimensional slices of the solution using each reconstruction methods: (a) $x=0.89$, (b) $x=0.5$.
Shock-vortex interaction: In this problem, we consider the interaction of a left-moving shock wave with a right-moving vortex [14]. The domain is $[0,1]\times[0,1]$, final time is $T=0.2$, and CFL = 0.45. The initial shock is given by
${\bm{u}}_{0}(x)=\begin{cases}{\bm{u}}^{L}&\text{if }x<0.5\\ {\bm{u}}^{R}&\text{if }x\geq 0.5\end{cases},$
where the left state is ${\bm{u}}^{L}=(\rho^{L},\text{u}_{1}^{L},\text{u}_{2}^{L},p^{L})=(1,\sqrt{\gamma},0,1)$ and the right state is given by
$\hskip 14.22636ptp^{R}$ $=1.3,\hskip 8.53581pt\rho^{R}=\rho^{L}\left(\frac{\gamma-1+(\gamma+1)p^{R}}{\gamma+1+(\gamma-1)p^{R}}\right),$
$\hskip 14.22636pt\text{u}_{1}^{R}$ $=\sqrt{\gamma}+\sqrt{2}\left(\frac{1-p^{R}}{\sqrt{\gamma-1+p^{R}(\gamma+1)}}\right),\hskip 8.53581pt\text{u}_{2}^{R}=0.$
A vortex is introduced into the left state by the following perturbations
$\hskip 14.22636pt\delta\text{u}_{1}$ $=\epsilon\frac{(y-y_{c})}{r_{c}}\exp{\left(\kappa(1-r^{2})\right)},\hskip 8.53581pt\delta\text{u}_{2}=-\epsilon\frac{(x-x_{c})}{r_{c}}\exp{\left(\kappa(1-r^{2})\right)},$
$\hskip 14.22636pt\delta\mathcal{T}$ $=-\frac{\gamma-1}{4\kappa\gamma}\epsilon^{2}\exp{\left(2\kappa(1-r^{2})\right)},\hskip 8.53581pt\delta s=0.$
where $r^{2}=((x-x_{c})^{2}+(y-y_{c})^{2})/r_{c}^{2}$. Note that the perturbation represents a steady-state isentropic vortex, whose parameters are selected to be $\epsilon=0.3$, $r_{c}=0.05$, $\kappa=0.204$, and $(x_{c},y_{c})=(0.25,0.5)$. The boundary conditions are Neumann, and the mesh consists of 400 $\times$ 400 cells. From the solutions shown in Figure 16, all methods seem to perform equally well for this test case.
@@
Figure 16: Shock-vortex interaction: Density profiles at time $T=0.2$. Comparison of TeCNO4 with different reconstruction methods: (a) ENO3, (b) SP-WENO, (c) SP-WENOc, (d) DSP-WENO.
6 Conclusion
In this work, we designed a novel neural network-based third-order WENO scheme called DSP-WENO, which is guaranteed to satisfy the sign property. Thus, DSP-WENO can be used within the TeCNO framework to obtain high-order entropy stable finite difference schemes. The motivation behind the proposed approach was two-fold: i) overcome the linear instability issues faced by ENO reconstruction, and ii) have better shock-capturing capabilities compared to existing WENO algorithms, i.e., SP-WENO and SP-WENOc, that satisfy the sign property.
A key element in the proposed strategy was to decouple the constraints guaranteeing the sign property and third-order accuracy (in smooth regions) from the learning process. A strong imposition of these constraints led to a convex polygonal feasible region from which the WENO weights need to be selected. Then a network was trained to adaptively choose the weights from the feasible region to ensure good reconstruction properties near discontinuities. In contrast, we could impose these constraints weakly by adding a penalty term to the loss functions (analogous to physics informed neural networks [43]). However, this would lead to the following challenges:
The training data used to train DSP-WENO did not require solving conservation laws. Instead the data was generated from a library of functions with varying smoothness, which mimic the local structures typically arising in solutions to conservation laws. Thus, the cost of generating training data is insignificant, and the trained DSP-WENO is model agnostic, i.e., does not depend on a specific conservation law. In other words, the DSP-WENO needs to be trained once offline and can then be used for any conservation law. This strategy was based on similar ideas first considered in [45] for designing troubled-cell detectors.
When comparing the numerical solutions obtained using various reconstruction methods satisfying the sign property, DSP-WENO achieved third-order accuracy in smooth regions, while being more dissipative compared to SP-WENO and SP-WENOc. This was attributed to the fact that the reconstructed jump is mostly zero (or small) with SP-WENO and SP-WENOc. However, this had a negative impact near discontinuities where the solutions exhibited large spurious oscillations due to insufficient diffusion. DSP-WENO significantly mitigated these spurious oscillations without compromising the order of accuracy in smooth regions. Further, DSP-WENO did not suffer from linear instabilities faced by ENO3. One could argue that ENO3 results in marginally better solutions near discontinuities compared to DSP-WENO. However, it is important to acknowledge that the ENO3 stencil (corresponding to an interface) comprises six cells, while SP-WENO and its variants (including DSP-WENO) have access to the solution on a stencil with just four cells.
The present work demonstrates that it is possible to use deep learning tools to learn adaptive reconstruction algorithms constrained to satisfy critical physical properties, such as the sign property. Further, it provides a framework to extend DSP-WENO to high-order ($>3$) accurate reconstructions satisfying the sign property. This would involve formulating the sign property constraint on a larger stencil and transforming the constraints into corresponding convex polyhedral regions in high dimensional spaces for the WENO weights. Thus, instead of attempting to construct an explicit weight selection strategy by hand (which presents significant challenges in high dimensions), a neural network can learn the selection procedure from training data. Further, similar sign-preserving WENO-type reconstructions can also be designed on unstructured grids. These extensions will be explored in future work.
Acknowledgements

### v2

@@ 1 $\theta_{i}^{+},\theta_{i+1}^{-}>1$ $C_{1}=C_{2}=\frac{1}{8}$ None $\tilde{w}_{0}=w_{1}=0$
Figure SM2: Reconstruction of a simulated discontinuity, which is not well-resolved, using: (a) ENO3, (b) SP-WENO, (c) SP-WENOc, (d) DSP-WENO.
SM4.2 Linear Advection
Tables SM2 and SM3 show the errors (measured in the discrete $L^{2}$ and $L^{\infty}$ norm, respectively) with various reconstructions for both smooth linear advection test cases. As with $L^{1}$, we observe third-order accuracy in $L^{2}$ in both test cases (except for ENO 3 in Test case 2). There seems to be slight deterioration in the $L^{\infty}$-order of convergence (which is much more severe for ENO3 in Test case 2).
@@
Test 1 100 1.43e-5 - 6.08e-5 - 5.69e-5 - 5.77e-5 -
200 1.79e-6 3.00 8.40e-6 2.86 7.91e-6 2.85 5.79e-6 3.32
400 2.24e-7 3.00 1.17e-6 2.85 1.13e-6 2.81 6.89e-7 3.07
600 6.63e-8 3.00 3.64e-7 2.87 3.60e-7 2.82 2.05e-7 2.99
800 2.80e-8 3.00 1.59e-7 2.87 1.57e-7 2.88 8.74e-8 2.96
1000 1.43e-8 3.00 8.40e-8 2.86 8.33e-8 2.84 4.47e-8 3.01
Test 2 100 7.02e-4 - 9.88e-4 - 9.71e-4 - 1.14e-3 -
200 9.44e-5 2.90 1.35e-4 2.87 1.34e-4 2.86 1.37e-4 3.05
400 1.24e-5 2.93 1.83e-5 2.89 1.81e-5 2.89 1.68e-5 3.03
600 4.04e-6 2.76 5.78e-6 2.84 5.76e-6 2.82 5.38e-6 2.80
800 2.51e-6 1.66 2.50e-6 2.92 2.48e-6 2.94 2.49e-6 2.67
1000 2.15e-6 0.68 1.33e-6 2.83 1.31e-6 2.84 1.32e-6 2.85
Table SM2: $L^{2}$ errors and convergence rates for linear advection smooth tests 1 and 2.
N ENO3 SP-WENO SP-WENOc DSP-WENO
Error Rate Error Rate Error Rate Error Rate
Test 1 100 9.20e-6 - 1.05e-4 - 9.77e-5 - 1.13e-4 -
200 1.10e-6 3.06 2.02e-5 2.38 1.91e-5 2.36 1.11e-5 3.35
400 1.42e-7 2.96 3.56e-6 2.50 3.47e-6 2.46 1.71e-6 2.69
600 4.21e-8 2.99 1.24e-6 2.60 1.23e-6 2.56 5.94e-7 2.61
800 1.73e-8 3.10 6.42e-7 2.30 6.33e-7 2.31 2.98e-7 2.39
1000 8.88e-9 2.98 3.83e-7 2.31 3.80e-7 2.28 1.70e-7 2.51
Test 2 100 6.38e-4 - 1.33e-3 - 1.30e-3 - 1.75e-3 -
200 9.22e-5 2.79 2.21e-4 2.59 2.19e-4 2.57 2.07e-4 3.08
400 1.29e-5 2.84 4.16e-5 2.41 4.14e-5 2.40 2.82e-5 2.88
600 4.44e-6 2.62 1.45e-5 2.59 1.45e-5 2.59 1.15e-5 2.21
800 3.56e-6 0.76 7.62e-6 2.24 7.49e-6 2.29 6.08e-6 2.21
1000 3.24e-6 0.43 4.45e-6 2.41 4.37e-6 2.41 3.84e-6 2.06
Table SM3: $L^{\infty}$ errors and convergence rates for linear advection smooth tests 1 and 2.
SM4.3 2D Riemann problem (conf. 12)
Figure SM3 shows the density contour plots for the 2D Riemann problem (configuration 12) when solved using ENO3, SP-WENO, SP-WENOc, and DSP-WENO.
(a) ENO3
(b) SP-WENO
(c) SP-WENOc
(d) DSP-WENO
@@
Figure SM3: 2D Riemann problem (configuration 12): Density profiles at time $T=0.25$ with 30 contour lines between 0.52 and 2.2. Comparison of different reconstruction methods
SM4.4 2D Riemann problem (conf. 3)
Figure SM4 shows the density contour plots for the 2D Riemann problem (configuration 3) with Rusanov dissipation when using ENO3, SP-WENO, SP-WENOc, and DSP-WENO. Figure SM5 shows a zoomed-in perspective of the carbuncle-like artifact for SP-WENO and DSP-WENO with the Roe and Rusanov dissipation. We observe that switching to the Rusanov dissipation mitigates the behavior.
(a) ENO3
(b) SP-WENO
(c) SP-WENOc
(d) DSP-WENO
@@
Figure SM4: 2D Riemann problem (configuration 3): Density profiles at time $T=0.3$ with 30 contour lines between 0.1 and 2.28. Comparison of different reconstruction methods
(a) SP-WENO Roe
(b) DSP-WENO Roe
(c) SP-WENO Rusanov
(d) DSP-WENO Rusanov
Figure SM5: 2D Riemann problem (configuration 3): Zoom-in of the carbuncle-like phenomenon for SP-WENO and DSP-WENO with Roe and Rusanov dissipation.
SM4.5 Kelvin-Helmholtz Instability
Figures SM6 and SM7 show the density profiles for the Kelvin-Helmholtz instability when solved using ENO3, SP-WENO, SP-WENOc, and DSP-WENO for meshes consisting of 256 $\times$ 256 cells and 512 $\times$ 512 cells.
(a) ENO3
(b) SP-WENO
(c) SP-WENOc
(d) DSP-WENO
@@
Figure SM6: Kelvin-Helmholtz Instability: Density profiles at time $T=3$. Comparison of different reconstruction methods for $256\times 256$ mesh.
(a) ENO3
(b) SP-WENO
(c) SP-WENOc
(d) DSP-WENO
Figure SM7: Kelvin-Helmholtz Instability: Density profiles at time $T=3$. Comparison of different reconstruction methods for $512\times 512$ mesh.